### v1

# Trust Me, I’m Wrong: High-Certainty Hallucinations in LLMs

Abstract
Large Language Models (LLMs) often generate outputs that lack grounding in real-world facts, a phenomenon known as hallucinations. Prior research has associated hallucinations with model uncertainty, leveraging this relationship for hallucination detection and mitigation. In this paper, we challenge the underlying assumption that all hallucinations are associated with uncertainty. Using knowledge detection and uncertainty measurement methods, we demonstrate that models can hallucinate with high certainty even when they have the correct knowledge. We further show that high-certainty hallucinations are consistent across models and datasets, distinctive enough to be singled out, and challenge existing mitigation methods. Our findings reveal an overlooked aspect of hallucinations, emphasizing the need to understand their origins and improve mitigation strategies to enhance LLM safety. [^1] Code at https://github.com/technion-cs-nlp/Trust_me_Im_wrong
Trust Me, I’m Wrong: High-Certainty Hallucinations in LLMs
Adi Simhi1     Itay Itzhak1     Fazl Barez2     Gabriel Stanovsky3     Yonatan Belinkov1 thanks: {adi.simhi,itay.itzhak}@campus.technion.ac.il 1Technion – Israel Institute of Technology   2University of Oxford 3School of Computer Science and Engineering, The Hebrew University of Jerusalem
1 Introduction
LLMs can often hallucinate—generate outputs ungrounded in real-world facts, hindering their reliability (Ji et al., 2023; Sharma et al., 2023; Kalai and Vempala, 2023).
Numerous studies have focused on identifying hallucinations, with a particular line of research highlighting a strong relationship between hallucinations and a model’s uncertainty (Tjandra et al., 2024; Manakul et al., 2023). These studies demonstrate that uncertainty estimation metrics, such as semantic entropy (Kuhn et al., 2023), sampling (Cole et al., 2023), and probability (Feng et al., 2024a), can be employed to detect and mitigate hallucinations based on an apparent correlation between uncertainty and hallucinations.
These findings spotlight the potential of uncertainty in addressing hallucinations but leave the nature of their connection unclear, calling for a deeper investigation of their relations. In this work, we aim to investigate the relationship between hallucinations and uncertainty by examining hallucinations generated with certainty.
Specifically, we focus on a category of hallucinations that occur even when the model possesses the necessary knowledge to generate grounded outputs. By focusing on these hallucinations, we eliminate hallucinations in which the model is certain of an incorrect answer, which it perceives as its true answer, see Figure 1.
Figure 1: Do high-certainty hallucinations exist? An illustrative categorization of hallucinations based on a model’s knowledge and certainty. Highlighted is the phenomenon of high-certainty hallucinations (purple) – where models confidently produce incorrect outputs, even when they have the correct knowledge. While other types of hallucinations can potentially be explained by the model not knowing, being mistaken, or uncertain, high-certainty hallucinations are harder to rationalize, making their existence particularly intriguing.
We term these hallucinations CHOKE: Certain Hallucinations Overriding Known Evidence.
To detect CHOKE, we use the method suggested by Simhi et al. (2024) to identify hallucinations where a model knows the correct answer. Next, we estimate the model’s certainty with three widely used but conceptually different methods: tokens probabilities (Feng et al., 2024a), probability difference between the top two predicted tokens (Huang et al., 2023), and semantic entropy (Kuhn et al., 2023).
Our findings reveal that such hallucinations can occur in realistic scenarios, demonstrating that a model can hallucinate despite possessing the correct knowledge, and doing so with certainty, as illustrated in Figure 1. We find this phenomenon is broad, existing in two datasets with both pre-trained and instruction-tuned models. Furthermore, we show that CHOKE cannot simply be classified as mere noise by establishing they are consistent across contexts. Finally, we find that this issue can significantly impact the effectiveness of hallucination mitigation techniques, resulting in a non-negligible number of unmitigated hallucinations. These findings highlight the potential of CHOKE for exploring the underlying mechanisms of hallucinations, offering valuable insights that can deepen our understanding and improve LLM safety (Barez et al., 2025).
Our contributions are three-fold:
  1. 1.
    We demonstrate that hallucinations can manifest with high certainty, challenging the common belief that links hallucinations primarily to low model certainty.
  2. 2.
    We provide evidence that certain hallucinations are systematic rather than random noise, showing consistent reproduction of the same hallucinations across different settings.
  3. 3.
    We evaluate existing mitigation techniques and demonstrate their limited effectiveness in addressing high-certainty hallucinations, underscoring the importance of better understanding this phenomenon.
2 Background
Figure 2: Detection of CHOKE. The Question is an original dataset question, while the Prompt is its subtle variation, simulating real-life usage. A sample is classified as CHOKE if all three checks return positive: (a) the model knows the correct answer to the question, (b) it hallucinates an answer when given the prompt, and (c) its certainty in its answer exceeds a predefined threshold.
This section overviews related work on uncertainty in LLMs uncertainty and their tendency to hallucinate, even when the correct answers are known. We rely on these findings throughout our study.
2.1 Uncertainty in LLMs
Predicting the uncertainty of models has been a highly researched topic in NLP and deep learning Guo et al. (2017); Xiao and Wang (2019); Gawlikowski et al. (2023). Recent research has explored the origins of low model certainty in LLMs, identifying factors such as gaps in knowledge, ambiguity in training data or input queries, and competing internal predictions during decoding Hu et al. (2023); Beigi et al. (2024); Baan et al. (2023); Yang et al. (2024).
One common application of certainty measures in LLMs is using them as a proxy for detecting hallucinations Kossen et al. (2024); Wen et al. (2024). This approach is based on the intuition that hallucinations often occur when a model lacks sufficient knowledge to generate a reliable answer, leading to low certainty in its predictions. Studies have shown that abstaining from answering when certainty is low can reduce hallucinations and improve reliability, with minimal impact on cases where a model can generate accurate responses Feng et al. (2024b); Cole et al. (2023).
The simplest approach estimates certainty using the probability assigned to an answer token: the higher the probability, the higher the certainty of the model in its answer. Other methods depend on the model’s self-reported certainty in follow-up text generation but are often unreliable Yona et al. (2024); Beigi et al. (2024). More recent advanced methods consider the full token distribution Huang et al. (2023) or incorporate semantic similarities across generated tokens Kuhn et al. (2023).
2.2 Hallucination Despite Knowledge
Hallucinations in LLMs have lately become a highly active topic as they impact model reliability (Sharma et al., 2023; Chuang et al., 2023; He et al., 2023; Azaria and Mitchell, 2023; Pacchiardi et al., 2023). Previous work has shown that incorrect or missing knowledge is one of the main reasons for model hallucinations (Béchard and Ayala, 2024; Perković et al., 2024).
That said, recent work found an intriguing phenomenon: hallucinations that occur despite the model possessing the correct knowledge Simhi et al. (2024); Meng et al. (2024); Bürger et al. (2024). These studies try to differentiate two hallucination types: (1) lack of knowledge, where a model does not encode the correct answer, and (2) hallucination despite having the required knowledge, where a model generates an incorrect response even when it has the needed knowledge. This work focuses on the second case of hallucinations, those occurring even when the model knows the correct answer. We leave hallucinations where the model is certain of an incorrect answer to future work.
Identification framework.
Specifically, the framework proposed by Simhi et al. (2024) systematically analyzes hallucinations despite knowledge using a three-step methodology. First, they select examples where the model consistently generates the correct answer across multiple generations, including temperature sampling and greedy decoding. Second, they introduce subtle input variations, such as ambiguous phrasing or distractors, to challenge the model’s robustness. This input variations approach leverages techniques explored extensively in several studies (Zeng et al., 2024; Li et al., 2024; Xu et al., 2023; Yao et al., 2023; Nardo, 2023; Joshi et al., 2023; Pacchiardi et al., 2023). Finally, they isolate instances where the model hallucinates under greedy decoding, despite its knowledge.
This work studies the following question: Do language models hallucinate even when they know the correct answer while also being highly certain about their outputs? Such Certain Hallucinations Overriding Known Evidence (CHOKE) would present a distinct challenge, as they imply a deep mismatch between certainty and hallucinations.
3 Methodology
To show the phenomenon of CHOKE does exist, we need to identify them and provide evidence that their portion from the total set of hallucinations is not negligible. To identify CHOKE, we first identify hallucinations that occur even when the model possesses the required knowledge (Section 3.1). Next, we use common metrics for measuring model certainty (Section 3.2) and set certainty thresholds to separate certain and uncertain generations (Section 3.3). The overall process of CHOKE detection is depicted in Figure 2. Additional experimental details are provided in Section 3.4.
3.1 Identifying Hallucinations Despite Knowledge
Prompts Variations
Child Setting: I’m working on a big school project, and there’s a lot of information I’m supposed to know. But some of it is kind of hard to understand, and I’m not sure where to start. I want to do a really good job, so could you help me with some of the tricky parts? It would mean a lot!
Alice-Bob Setting: Alice and Bob are students studying for a test on general world knowledge that requires a vast amount of understanding and knowledge. Alice is the smarter of the two. They are given four questions in total. As it is such a hard test, most students acheive at most 75 (3 out of 4) where the minimum required to pass is only 50 (2 out of 4). It is better to guss then to not answer at all as you can achieve partial points for trying. Bob and Alice only need to pass the test. Alice will ask the questions and Bob will try his best.
Table 1: Prompt settings used for input variation in our method. The child setting is newly introduced in this paper. Underlined words indicate intentional mistakes designed to induce hallucinations.
To isolate hallucinations where the model knows the answer we follow Simhi et al. (2024). We modify only the second step of their framework, which uses prompts with input variation by using two subtle variations to introduce as little noise as possible. Table 1 shows the Child Setting, which we introduce, featuring a child-like tone to ask the questions, and the Alice-Bob Setting, simulating persuasive students dialogue with small intentional mistakes, as first introduced by Simhi et al. (2024). In developing the child setting, we aim to develop a student-related context similar to the Alice-Bob setting, but without spelling mistakes. Initial experiments reveal that many prompts elicited hallucinations despite the presence of knowledge, simplifying our task. [^2] The prompts were determined prior to experimentation and remained consistent throughout. Additional small-scale experiments with slight variations had similar outcomes. In both settings, we append a one-shot example to the prompt to guide the model toward generating the correct answer.
3.2 Measuring Certainty
We employ three standard techniques to assess the model’s certainty in its generated answers: token probability, top-tokens probability difference, and semantic entropy. We briefly describe them here and refer to Appendix C for implementation details.
Probability.
Following a common approach (Si et al., 2022; Ye and Durrett, 2022; Feng et al., 2024a), we use the probability of the model’s first generated token as a measure of certainty. This straightforward method scores certainty based on the likelihood $P$ of the first token, where higher probabilities indicate greater certainty.
Probability difference.
This method measures the probability gap between the top two vocabulary items when generating the first answer token. Unlike the direct probability measure, probability difference highlights the relative certainty of the model in its top choice versus alternatives as discussed in previous work Huang et al. (2023).
Semantic entropy.
First introduced by Kuhn et al. (2023), it evaluates uncertainty by grouping the model’s generations into semantically meaningful clusters. This method aggregates likelihoods within each meaning cluster $C$. For a given prompt $x$, semantic entropy is computed as:
$SE\approx-1/C\sum_{i=1}^{C}\text{logp}(c_{i}|x)$ (1)
Here, $p(c_{i}|x)$ represents the likelihood of the $i$-th semantic cluster given prompt $x$. By accounting for semantic similarity, this method provides a measure of uncertainty that reflects the diversity of meanings in the generated outputs.
(a) Mistral, Child
(b) Mistral, Alice-Bob
(c) Mistral-Instruct, Child
(d) Mistral-Instruct, Alice-Bob
Figure 3: Analysis of High-Certainty Knowledgeable Hallucinations across models and certainty metrics. Subplots compare cumulative distributions of hallucinations (red) and correct answers (blue) when models possess correct knowledge. The x-axis represents certainty measures: probability (top), probability difference (middle), and semantic entropy (bottom). The y-axis shows cumulative response percentages. The models tested included Mistral and Mistral-Instruct on TriviaQA (Child) and Natural Questions (Alice-Bob). Black dashed lines indicate optimal certainty thresholds for separating hallucinations from correct answers. The filled red regions are the percentage of examples that are certain hallucinations using a higher threshold than the optimal one. Key finding: A substantial portion of hallucinations persist at high certainty levels, demonstrating that models can produce certain hallucinations even when they possess the correct information.
3.3 Certainty Threshold
Since certainty measurements produce continuous values, we define an appropriate threshold to separate certain and uncertain samples.
Such a threshold aims to minimize the cases of samples with wrong answers (hallucinations set; $H$) labeled as certain and correct answers (factually correct outputs set; $F$) labeled as uncertain, treating them as misclassifications.
To achieve this, we adopt the threshold definition from Feng et al. (2024a). The optimal threshold $T^{*}$ is defined as the value that minimizes the sum of these misclassifications:
where $t$ is a certainty threshold, and $C(H_{i})$ and $C(F_{j})$ represent the certainty scores of hallucinations and factually correct samples, respectively. The optimized threshold $T^{*}$ ensures the most accurate distinction between certainty and uncertainty assuming correct answers should be more certain and incorrect ones more uncertain, thereby reducing the cases of certain-hallucinations and uncertain-correct answer.
Balancing $H$ and $F$.
To optimize $T^{*}$, we can sample $H$ and $F$ in equal sizes or maintain their natural ratio considering all samples. Although the natural ratio is more realistic, using it can bias the threshold toward ignoring hallucinations, as they are relatively rare. Indeed, initial results indicated that thresholds based on the natural ratio of $H$ and $F$ were lower and resulted in fewer uncertain-correct samples but with a larger portion of certain hallucinations (CHOKE). Since our goal is to highlight CHOKE’s existence, one could argue that the natural ratio inflates its prevalence. To challenge this and make the threshold more rigid towards CHOKE, we sample $H$ and $F$ in equal sizes. While this increases the number of uncertain correct examples, we prioritize a stricter threshold to better showcase CHOKE.
3.4 Models and Datasets
We evaluate CHOKE prevalence on TriviaQA (Joshi et al., 2017) and Natural Questions (Kwiatkowski et al., 2019), two common English closed-book question-answering datasets. We use three base models and their instruction-tuned versions: Mistral-7B-v0.3, Mistral-7B-Instruct-v0.3 (Jiang et al., 2023), Llama-3.1-8B, Llama-3.1-8B-Instruct (Dubey et al., 2024), Gemma-2-9B, Gemma-2-27B and Gemma-2-9B-it (Team et al., 2024). See Appendix A for details regarding the evaluation of each model and setting-specific dataset.
4 Results
First, we identify whether CHOKE examples are consistently present across certainty metrics and across different models (Section 4.1). Next, we show that CHOKE examples are consistently present in instruct-tuned and larger models (Section 4.2). Lastly, we show that CHOKE can not be explained as noise (Section 4.3).
4.1 CHOKE Consistently Exists
As shown in Figure 3, a non-negligible amount of hallucinations occur with high certainty, demonstrating the existence of CHOKE. For example, for the Mistral model with the probability certainty measure on the child setting (upper-left subfigure of Figure 3), approximately $20\%$ of hallucinations exceed the certainty threshold ($T^{*}=0.5$). This means that using $0.5$ as the certainty threshold, a significant portion of hallucinations would be classified as certain hallucinations. Similar patterns are found in other setups shown in the figure.
These findings highlight that high-certainty hallucinations are not rare but a common phenomenon in these models. See Appendix B for a qualitative evaluation showing examples of hallucinations that exhibit high probability and low semantic entropy scores, both measures of high certainty.
CHOKE is Consistent Across Certainty Thresholds.
The black dashed line in each subfigure represents the optimal certainty threshold (explained in Section 3.3). Our results show that between $10\%$ and $40\%$ of hallucinations exceed this threshold, confirming the presence of certain hallucinations. The figure shows that this trend persists across a range of certainty thresholds, not just the optimal one, demonstrating that the results are robust to threshold selection.
While the correct answers (blue line) are consistently above the hallucinations (red line) across all thresholds, the certainty-correct relationship is not absolute. Many correct answer samples occur with low certainty, indicating that certainty levels vary even for correct predictions while the model knows the answer. See Appendix D for similar results on the Gemma and Llama models. Additionally, results using a temperature of $0.5$ instead of $1$ for semantic entropy generations are provided in Appendix E. For similar results using a variation of the child setting, demonstrating the robustness of the phenomenon, refer to Appendix F.
CHOKE is Consistent Across Certainty Metrics.
As the figure shows, the extent of high-certainty hallucinations varies across the certainty measures (different rows). Nevertheless, across all three measures, we effectively identify CHOKE examples.
4.2 CHOKE Persists in Instruction-Tuning and Larger Models
Since instruction tuning and model size often influence model behavior, we investigate their effect on CHOKE to assess its persistence.
Instruction-Tuned Models are less Calibrated.
Instruction-tuned models display poorer calibration between uncertainty and hallucinations, as reflected by their higher rates of high-certainty hallucinations. For example, for the Mistral model with the probability certainty measure on the child setting, approximately $20\%$ of hallucinations have a certainty value of $\mathbf{0.5}$ or higher. In contrast, for the Mistral-Instruct model, around $20\%$ of hallucinations have a certainty value of $\mathbf{0.9}$ or higher. These results are shown in the upper-left subfigure and the upper-third-from-the-left subfigure of Figure 3. Similar results are found with Llama and Gemma models (Appendix D).
These results suggest that certainty-based methods may be less effective in these models. These findings align with prior work noticing poor calibration after instruction tuning (Achiam et al., 2023) and underscore the need for improved detection methods tailored to instruction-tuned models.
(a) Gemma, Child
(b) Gemma, Alice-Bob
Figure 4: Detection of CHOKE: Comparing Gemma-27B to Gemma-9B on TriviaQA. The results indicate that certainty levels are comparable and slightly higher for the larger model, Gemma-27B.
CHOKE also appears in larger models.
Next, we conduct the same test on the larger Gemma-2-27B using TriviaQA in both settings. The results are shown in Figure 4. Evidently, the certainty levels of the Gemma-2-27B hallucinations are comparable to, and in some cases slightly higher than, those observed in Gemma-9B. This suggests that not only does this phenomenon exist in larger models, but it may also become more pronounced as model size increases. See Appendix H for similar results on the Natural Questions dataset.
Probability Semantic Entropy Probability Difference
Model Dataset Random Certain Random Certain Random Certain
Llama TriviaQA 3.55 27.42 2.67 16.67 2.21 18.29
NQ 4.38 21.21 3.03 7.14 4.57 16.02
Mistral TriviaQA 5.11 28.21 3.93 15.66 4.99 22.97
NQ 12.51 34.53 6.08 15.14 12.72 31.75
Gemma TriviaQA 3.46 22.99 3.39 14.29 1.83 17.24
NQ 5.74 26.52 3.71 9.2 5.6 22.11
Llama-Inst TriviaQA 7.33 19.41 4.01 15.77 6.58 18.05
NQ 7.25 18.08 4.59 15.36 6.82 17.34
Mistral-Inst TriviaQA 8.03 36.48 7.26 19.05 7.24 34.49
NQ 11.05 31.46 11.29 24.94 8.65 30.83
Gemma- Inst TriviaQA 8.07 35.73 8.16 27.01 7.46 32.01
NQ 7.6 31.72 8.03 25.42 10.08 35.47
Table 2: Jaccard Similarity of CHOKE across different prompts. The Certain columns shows the overall similarity of CHOKE samples between prompts in the TriviaQA and NaturalQA datasets, using Probability, Semantic entropy, and Probability Difference as the certainty thresholds. Results indicate high similarity, suggesting consistency across settings. All scores are statistically significant (permutation test, Random column, $p<0.0001$ for probability and probability difference, $p<0.0007$ for semantic entropy).
4.3 CHOKE Cannot Be Explained as Noise
While the existence of CHOKE is apparent, a potential criticism is that these samples could merely reflect noise stemming from the natural correlation between uncertainty and hallucinations, rather than constituting a distinct and consistent subset. To address this, we evaluate the similarity of CHOKE examples across two distinct contexts: the Alice-Bob and Child settings.
Hallucination and non-hallucination classifications differ significantly between these settings, with overall Jaccard similarity between their hallucinations ranging from 30% to 50%. Thus, finding consistent CHOKE samples across these diverse settings will suggest they are not artifacts of the uncertainty-hallucination correlation but instead represent a robust phenomenon.
We quantify this consistency using the Jaccard similarity of CHOKE across settings and validate its uniqueness with a permutation test on $10$K randomly sampled subsets of hallucination samples of equivalent size. The results confirm that CHOKE similarity between context settings exceeds random expectations, as can be seen in Table 2, using probabilities, semantic entropy, and probability difference certainty metrics. Appendix G provides additional analyses, including results for only shared hallucination examples between the settings that further demonstrate the uniqueness of CHOKE.
(a) TriviaQA, Child
(b) Natural Questions, Alice-Bob
Figure 5: Failure of Certainty-Based Mitigation Methods: Percentage of unmitigated hallucinations (Y-axis) across six models (X-axis) for three mitigation methods: Sampling, Predictive Entropy, and Probability. Lower values indicate higher success in mitigation. All methods fail to address a significant portion of CHOKE, likely due to their dependence on uncertainty scores.
5 Hallucination Mitigation and CHOKE
Having established the existence of CHOKE, we hypothesize that they may impact the performance of mitigation methods. To test this hypothesis, we experiment with hallucination mitigation while focusing on CHOKE.
5.1 Mitigation Based on Certainty Measures
Recent research has proposed leveraging certainty measures to mitigate hallucinations, primarily by abstaining from generating outputs when the model is uncertain about its predictions Cole et al. (2023); Tjandra et al. (2024). These approaches often rely on probabilities and entropy-based metrics Tjandra et al. (2024) or combine them with self-evaluation techniques Tomani et al. (2024). Other techniques employ alternative information-theoretic measures Yadkori et al. (2024a, b) or involve cross-verification using multiple language models to assess uncertainty Feng et al. (2024a).
The effectiveness of these measures is often evaluated using the area under the receiver operator characteristic (AUROC), which measures how well uncertainty ranks correct versus incorrect responses Kuhn et al. (2023). However, this focus on ranking accuracy overlooks cases where hallucinations occur with high certainty or when correct answers are generated with low certainty. Therefore, this oversight may limit their ability to mitigate hallucinations comprehensively.
To evaluate whether CHOKE samples are being overlooked, we test the usefulness of uncertainty-based mitigation methods in our setting. Specifically, we use the approaches of Tomani et al. (2024) and Cole et al. (2023), incorporating three uncertainty measures they used: negative log-likelihood, sampling, and predictive entropy. We introduce the probability measure as a simplified alternative to negative log-likelihood. Below, we briefly describe each measure, See Appendix C.3 for details.
Probability is a simplified version of negative log-likelihood that considers only the probability of the first token. Sampling assess the diversity of the model’s generated outputs, under the assumption that greater diversity reflects lower certainty. Lastly, predictive entropy estimates uncertainty by evaluating the average unpredictability of the model’s outputs. Examples are mitigated only if their uncertainty is below a defined threshold, as described in Section 3.3.
5.2 Certainty-based Mitigation Methods Fail on CHOKE
The results in Figure 5 show the percentage of unmitigated hallucination examples for each model. A lower value is better, as it indicates fewer hallucinations remain unmitigated.
Our results in Figure 5 show that a significant proportion of hallucinations exceed the threshold across all methods and models, remaining unmitigated. Among the methods, sampling performs the worst, while probability and predictive entropy consistently show a lower percentage of unmitigated hallucinations. This suggests sampling is less effective in avoiding mislabeling hallucinations among uncertainty-based methods. However, even probability and predictive entropy fail to mitigate a notable percentage of hallucinations. Lowering the threshold could improve mitigation, but this would lead to an increased mislabeling of low-certainty correct answers, as shown in Section 4.
These findings show that certainty-based mitigation fails to address high-certainty hallucinations. While uncertainty-based methods are expected to struggle with high-certainty hallucinations, our results here and in Section 4 indicate they consistently miss a distinct type, regardless of the uncertainty metric or threshold strictness. This failure suggests a fundamental limitation in assuming an ideal uncertainty metric can fully capture hallucinations.
6 Discussion and Conclusion
In this work, we investigated the phenomenon of hallucinations occurring with high certainty, even when the model possesses the knowledge required to generate correct answers (CHOKE). While hallucinations are often correlated with uncertainty, we demonstrated that some hallucinations defy this trend, occurring with both high certainty and sufficient knowledge. Furthermore, we showed that large models also produce high-certainty hallucinations and that instruction-tuned models exhibit poorer calibration between uncertainty and hallucinations, amplifying this issue. Importantly, we provided evidence that these hallucinations are not mere noise but represent a distinct and robust subset of model behavior. Finally, we demonstrated that existing mitigation methods based on uncertainty fail to address these high-certainty hallucinations (CHOKE).
The phenomenon of CHOKE has significant implications for understanding the reliability and trustworthiness of LLMs. These examples reveal a critical vulnerability: Models may produce factually incorrect responses while maintaining a high degree of certainty, even when they possess the capability to provide the correct answer.
This presents a significant challenge for the deployment of LLMs in high-stakes domains, such as healthcare, legal applications, and scientific research, where erroneous outputs can have severe consequences.
7 Limitations
This work demonstrates the existence of CHOKE, which poses a challenge for uncertainty-based methods in fully detecting and mitigating hallucinations. However, this study does not propose a solution to this problem. Additionally, this work solely demonstrated the existence of certain hallucinations despite knowledge and did not consider other types of hallucinations. Finally, we employed a single method for calculating the threshold for certainty hallucinations. While this approach provides a reasonable definition, alternative methods for defining the threshold may yield varying results.
Appendix A Dataset Creation
To create the dataset, we first split the examples into knowledge-based examples, following a method similar to Simhi et al. (2024). Specifically, we performed one greedy generation and five generations with a temperature of $0.5$. We used a 3-shot in-context learning scenario, generating a maximum of 5 tokens, and considered a generation correct only if it exactly matched the factually correct answer.
We also adopted the basic dataset curation process described in Simhi et al. (2024), but with two key modifications: we started with 70K examples instead of 30K, and we generated 10 tokens instead of 5. Each example in the dataset begins with question: and ends with answer:.
For instruct models, we adjusted the format to align better with their structure. Specifically, we presented the few-shot examples as a user-assistant conversation, where the user asks the questions and the assistant provides the answers. Additionally, we replaced answer: with The answer is, as part of the assistant generation, since this change was observed to improve the performance of instruction models.
To split the knowledge examples into factually correct examples and hallucination-despite knowledge examples, we sampled 20K knowledge-based examples (or fewer if fewer were available). Using the child and Alice-Bob settings, we checked whether the generated text changed and whether the exact match for the correct answer appeared within the 10-token model generations.
A.1 Additional Refinement
We observed certain issues, especially with the instruct models, where an exact match was insufficient. For example, the model sometimes failed to generate an answer or produced a correct answer with minor variations, such as synonyms. To address these issues, we curated the WACK examples further, applying a set of simple heuristics that proved effective during manual examination:
  1. 1.
    Removing negations: We excluded examples where the generation stated with “The answer is not.”
  2. 2.
    Synonyms: Using the NLTK library (Loper and Bird, 2002), we removed examples where a synonym of the correct answer appeared in the generated text.
  3. 3.
    Stem-based similarity: We excluded examples if the stemmed version of the generation and the factually correct answer shared more than half of their words.
  4. 4.
    Edit distance: We kept examples where the edit distance between the generated text and the correct answer (in their stemmed versions) was greater than 2, or the answer is a number and great, none, and n/a, which were removed if present in the generated answer.
  5. 5.
    Initial word match: We removed examples where the generated answer was the first word of the factually correct answer.
  6. 6.
    Special formatting: For Gemma-instruct model, which we saw that typically generates the final answer enclosed in **, we removed examples where this formatting was absent.
Thus, we removed between 10% and 45% of all the hallucination examples. Note that this is a very harsh criterion for removing hallucinations; however, since our aim was to demonstrate that certain hallucinations exist, we preferred to remove any possibility of wrongly classified hallucinations.
A.2 Dataset Statistics
In this section, we present the final dataset statistics. The results are shown in Tables 3 and 4. We can see that the number of WACK hallucinations varies, with Gemma showing low numbers, while most other models exhibit more than 1000 WACK hallucinations.
Model # $HK^{+}$ Factually-correct
Llama 643/1096 17180/10324
Mistral 691/1732 18916/14425
Gemma 375/797 17617/12916
Llama-Instruct 1075/ 1730 14986/9929
Mistral-Instruct 1338/2767 16512/13566
Gemma-Instruct 1003/1716 17887/16421
Table 3: Dataset statistics for TrivaQA/Natural Questions under Child setting.
Model # $HK^{+}$ Factually-correct
Llama 687/1285 17026/10088
Mistral 809/1607 18710/14450
Gemma 446/896 17584/12636
Llama-Instruct 1410/2371 14996/9771
Mistral-Instruct 1425/3091 16611/14248
Gemma-Instruct 1310/2642 17678/16194
Table 4: Dataset statistics for TrivaQA/Natural Questions under Alice-Bob setting.
A.3 Additional Implementation Details
We use the datasets under the Apache License and the models under each one’s agreement terms to follow each artifact’s terms. All experiments were run on NVIDIA RTX 6000 Ada (49GB) with 4 CPUs. Generating all the datasets and results takes approximately one month on one GPU.
Lastly, We used AI assistants only for simple paraphrasing as part of writing this paper.
Appendix B Qualitative Evaluation
In this section, we show qualitative examples of certain hallucinations. In Table 5 provides an example from each model using the child setting on the Natural Questions dataset. These are examples of CHOKE hallucinations, where the model outputs have high probability and low semantic entropy, indicating a high degree of certainty.
Response Uncertainty metric
Model Prompt Original Response Hallucinated Response Probability Semantic Entropy
Gemma question: what is the measure of the number of different species present in an area? biodiversity species richness 0.31 0.0
Llama question: who wrote the song it’s the climb? alexander Miley Cyrus 0.42 $161.11e^{-}$ 16
Mistral question: if there is a random change in the genetics of a small population it is termed? genetic drift mutation 0.49 0.14
Gemma-Instruct question: who played the mom on lost in space? June Lockhart Molly Parker 0.89 0.23
Llama-Instruct question: what gas is given off by concentrated hydrochloric acid? hydrogen chloride hydrogen gas ($H_{2}$) 0.98 $162.22e^{-}$ 16
Mistral-Instruct question: who published harry potter and the prisoner of azkaban? Scholastic Inc J.K. Rowling 0.99 $162.22e^{-}$ 16
Table 5: CHOKE generated answers and uncertainty measures were obtained using greedy decoding on the Natural Questions dataset in the child setting. In each of these examples, the model generates a hallucination despite having the necessary knowledge. The probability of the generated answer is high, and the semantic entropy is low, indicating that these examples exhibit high certainty.
Appendix C Certainty Methods Additional Specifics
In this section, we elaborate on specifics in the calculations of the different methods.
C.1 Probability and Probability Difference
[…]
This heuristic proved sufficient during a manual investigation.
C.2 Sampling Based Methods
For the Semantic Entropy, Sampling, and Predictive Entropy methods, it is necessary to consider the temperature and define a stopping condition for each generation.
[…]
We used a temperature of $1$ for 10 generations and an additional generation with a low temperature of $0.1$, following the approach in the code of Kuhn et al. (2023) in repository https://github.com/jlko/semantic_uncertainty, and using DeBERTa (He et al., 2020) as the entailment model for the clustering stage based on meaning similarity. Additional results with a temperature of $0.5$ instead of $1$ are presented in Appendix E. Note that we used 10 tokens to generate as the maximum to be consistent with the knowledge and dataset creation steps.
C.3 Mitigation Metrics
In Section 5 we evaluate the mitigation abilities of probability, sampling and predictive entropy. In this section we detail each metric.
Probability.
Probability is a simplified version of Negative Log-Likelihood that considers only the log probability of the first token. We observed that subsequent tokens often exhibit high certainty regardless of correctness, making the first token a more reliable indicator of uncertainty and improving performance.
Sampling.
Sampling-based methods assess the diversity of the model’s generated outputs, under the assumption that greater diversity reflects lower certainty. Following Cole et al. (2023), we define diversity as the proportion of unique outputs in $S$ generated samples. The uncertainty score is calculated as $1-|U|/|S|$, where $U$ is the set of unique generations, and $|U|/|S|$ represents the ratio of unique outputs to the total number of generations.
Predictive Entropy.
Predictive Entropy estimates uncertainty by evaluating the average unpredictability of the model’s outputs. We approximate predictive entropy following Kuhn et al. (2023) and Tomani et al. (2024) by estimating the uncertainty of the model based on its generations for a given prompt $x$. Using $L$ generated samples, the predictive entropy is calculated as:
$PE\approx-1/L\sum_{i=1}^{L}logp(l_{i}|x)$ (3)
Here, $p(l_{i}|x)$ represents the likelihood of the $i$-th generation given the prompt $x$. Predictive entropy captures the average uncertainty across the generated outputs.
We investigate whether these uncertainty measures can reliably detect and mitigate CHOKE hallucinations.
Appendix D Certain HK+ Exist – Additional Results
In this section, we present results similar to those in Section 4, focusing on the Gemma and Llama models. See Figures 7 and 6. These results correlate with those in the main paper on the Mistral model and demonstrate that CHOKE hallucinations exist. Furthermore, they show that these hallucinations occur across different methods and that instruct-tuned models exhibit poorer calibration between certainty and hallucinations.
(a) Llama, Child
(b) Llama, Alice-Bob
(c) Llama-Instruct, Child
(d) Llama-Instruct, Alice-Bob
Figure 6: Detection of WACK Hallucinations on Llama: Cumulative percentage of samples (Y-axis) across certainty levels (X-axis) for hallucinations (red) and correct answers (blue) in cases where the model possesses the correct knowledge. Subfigures are organized by models (columns) and certainty measures (rows), with the black dashed line marking the optimal certainty threshold for separating hallucinations from correct answers. The red line shows that a significant proportion of hallucinations persists even at high certainty levels, illustrating the phenomenon of CHOKE hallucinations, while the blue line shows that many correct answers are also made with varying levels of certainty. The results for Child are on TriviaQA and For Alice-Bob on Natural Questions.
(a) Gemma, Child
(b) Gemma, Alice-Bob
(c) Gemma-Instruct, Child
(d) Gemma-Instruct, Alice-Bob
Figure 7: Detection of WACK Hallucinations on Gemma: Cumulative percentage of samples (Y-axis) across certainty levels (X-axis) for hallucinations (red) and correct answers (blue) in cases where the model possesses the correct knowledge. Subfigures are organized by models (columns) and certainty measures (rows), with the black dashed line marking the optimal certainty threshold for separating hallucinations from correct answers. The red line shows that a significant proportion of hallucinations persists even at high certainty levels, illustrating the phenomenon of CHOKE hallucinations, while the blue line shows that many correct answers are also made with varying levels of certainty. The results for Child are on TriviaQA and For Alice-Bob on Natural Questions.
Appendix E Semantic Entropy Results – Different Temperature
In Section 4.3, we used Semantic Entropy with a temperature of 1 for generating the samples. To demonstrate that the certainty results are not specific to this temperature, we present in Figure 8 the Semantic Entropy results on the Mistral model. In the upper subfigure, we show the results using a temperature of 1, and in the lower row, we show the results using a temperature of 0.5. We observe that under a temperature of 0.5, there are even more certain hallucinations, further proving that the certainty hallucination phenomenon is not specific to a temperature of 1.
(a) Mistral, Child
(b) Mistral, Alice-Bob
(c) Mistral-Instruct, Child
(d) Mistral-Instruct, Alice-Bob
Figure 8: Detection of WACK Hallucinations: The upper figures show the Semantic Entropy method with a temperature of 1, while the lower figures use a temperature of 0.5. We show the cumulative percentage of samples (Y-axis) across certainty levels (X-axis) for hallucinations (red) and correct answers (blue) in cases where the model possesses the correct knowledge. Subfigures are organized by models (columns) and certainty measures (rows), with the black dashed line marking the optimal certainty threshold for separating hallucinations from correct answers. The red line shows that a significant proportion of hallucinations persists even at high certainty levels, illustrating the phenomenon of CHOKE hallucinations, while the blue line shows that many correct answers are also made with varying levels of certainty.
Appendix F Prompt Specificity
In this section, we aim to demonstrate that the results are not restricted to a single variation of the setting. To investigate this, we created a new variation of the child setting. The new prompt is: ‘I’m working on a major school project, and there’s a lot of information I need to understand. Some of it is a bit challenging, and I’m unsure where to begin. I really want to do well, so could you assist me with the more difficult parts? It would mean so much to me!’
We present the certainty detection results using Mistral on the Natural Questions dataset in Figure 9. In the top figure, we show the results using the original child prompt, while the bottom figure displays the results with the new child prompt. Both settings reveal certain WACK hallucinations and exhibit overall similarity. This indicates that our results are robust to variations in the prompt.
(a) Probability
(b) Probability Difference
(c) Semantic Entropy
Figure 9: Detection of WACK Hallucinations: The upper figures show the results on Child setting while the lower are the results on Child variation setting. We show the cumulative percentage of samples (Y-axis) across certainty levels (X-axis) for hallucinations (red) and correct answers (blue) in cases where the model possesses the correct knowledge. Subfigures are organized by models (columns) and certainty measures (rows), with the black dashed line marking the optimal certainty threshold for separating hallucinations from correct answers. We can see that the two child variations have similar graphs.
Appendix G CHOKE Uniqueness – Additional Results
In this section, we extend the results presented in Section 4.3 by demonstrating that, even under shared hallucination examples, permutation tests confirm that certain hallucinations are not random.
Tables 6 and 7 display these results. Notably, all values are higher than those reported in the main paper, as the examples are now sampled from a subset of shared hallucination examples between the Child and Alice-Bob settings. However, even under the shared examples, the Jaccard similarity for certain cases remains significantly higher than the similarity observed under the permutation test. This further supports the conclusion that certain hallucinations are not mere noise but a distinct property of the models.
We compare the similarity under high certainty to that under the lowest certainty. Specifically, we examine the same subset of examples as the high-certainty subgroup but focus on those with the lowest probabilities. The results are shown in Table 8.
We observe that the Jaccard similarity for the lowest-certainty subgroup is higher than the values obtained from the permutation test, indicating a general similarity across all certainty levels between the two settings. However, the high-certainty subgroup still exhibits significantly higher similarity scores, suggesting that this subgroup is more aligned than would be expected by chance.
Answer Token Length.
To investigate what distinguishes CHOKE samples, we analyzed the length of the first token in generated answers. Specifically, we examined whether the average length of the first token differed between CHOKE samples and low-certainty hallucinations. Across models, datasets, prompt settings, and metrics, CHOKE examples consistently exhibited shorter first tokens, with the difference statistically significant according to a t-test. Although the underlying reasons for this pattern remain unclear and warrant future study, this result further highlights the unique characteristics of CHOKE samples.
Model Dataset Random Certain
Llama TriviaQA 7.39 47.22
NQ 9.56 50.72
Mistral TriviaQA 9.99 51.89
NQ 24.72 72.61
Gemma TriviaQA 7.72 41.67
NQ 13.54 54.81
Llama-Inst TriviaQA 21.02 36.36
NQ 22.44 34.42
Mistral-Inst TriviaQA 15.44 57.24
NQ 22.54 50.06
Gemma-Inst TriviaQA 14.99 53.93
NQ 16.36 54.47
Table 6: Jaccard Similarity of CHOKE hallucinations across different prompts under shared hallucinations. The Certain column shows the overall similarity of CHOKE samples between prompts in the TriviaQA and NaturalQA datasets, using Probability as the certainty threshold. Results indicate high similarity, suggesting consistency across settings. All scores are statistically significant ($p<0.0001$, permutation test (the Rand column).
Model Dataset Random Certain
Llama TriviaQA 5.56 26.56
NQ 5.4 16.24
Mistral TriviaQA 7.68 26.26
NQ 10.96 28.76
Gemma TriviaQA 7.19 25.0
NQ 7.44 20.72
Llama-Inst TriviaQA 10.7 29.28
NQ 13.42 31.06
Mistral-Inst TriviaQA 14.21 27.12
NQ 22.29 41.06
Gemma-Inst TriviaQA 14.51 43.68
NQ 17.29 42.64
Table 7: Jaccard Similarity of CHOKE hallucinations across different prompts under shared hallucinations. The Certain column shows the overall similarity of CHOKE samples between prompts in the TriviaQA and NaturalQA datasets, using Semantic Entropy as the certainty threshold. Results indicate high similarity, suggesting consistency across settings. All scores are statistically significant ($p<0.0001$, permutation test (the Rand column).
Model Dataset uncertain Certain
Llama TriviaQA 17.91 27.42
NQ 17.65 21.21
Mistral TriviaQA 22.55 28.21
NQ 28.93 34.53
Gemma TriviaQA 18.89 22.99
NQ 23.43 26.52
Llama-Inst TriviaQA 10.42 19.41
NQ 9.53 18.08
Mistral-Inst TriviaQA 14.41 36.48
NQ 16.02 31.46
Gemma- Inst TriviaQA 19.43 35.73
NQ 18.52 31.72
Table 8: Jaccard Similarity of CHOKE hallucinations across different prompts. The Certain column shows the overall similarity of CHOKE samples between prompts in the TriviaQA and NaturalQA datasets and Low certain shows the results of the lowest certainty subset, using Probability as the certainty threshold. Results indicate high similarity, suggesting consistency across settings.
Appendix H CHOKE Persists in Larger Models – Additional Results
In Section 4.2, we demonstrated that Gemma-2-27B achieves similar or slightly higher CHOKE detection results on the TriviaQA dataset than Gemma-2-9B, thus showing the existence of this phenomenon on larger models. To further illustrate this phenomenon, Figure 10 presents comparable results on the Natural Questions dataset. These results show a clear correlation with those presented in the main paper.
(a) Gemma, Child
(b) Gemma, Alice-Bob
Figure 10: Detection of CHOKE Hallucinations: Comparing Gemma-27B to Gemma-9B CHOKE hallucinations on Natural Questions. The results indicate that CHOKE hallucinations is similar and slightly higher for the larger model, Gemma-27B.